OpenStudy (anonymous):

6n + 11 = -7 + 3n for n.

OpenStudy (anonymous):

subtract 3n from both sides, then subtract 11 from both sides

OpenStudy (anonymous):



OpenStudy (imtiaz7):

6n +11= -7 +3n 6n-3n =-7-11 3n = -18 n= -18/3 n= -6